The density of copper at 293 K is 8,940 kg/m3 and its linear expansion coefficient is 17.0 x 10−6 K−1. Consider a hot cube of copper that is 1.0 cm on a side when its temperature is 1.356 K. What is the cube’s mass? 


1
Expert's answer
2022-05-15T17:02:57-0400

"m=\frac{\rho V}{1+3\alpha∆t}=8.5~g."